an + an-1 - 10an-2 + 8an-3


1
Expert's answer
2022-01-13T10:22:30-0500

Since, given question is incomplete, so we assume as follows:

Consider the recurrence relation "a_{n}=a_{n-1}+6 . a_{n-2}, \quad n \geq 2, a_{0}=3, a_{1}=6" The characteristic equation is "r^{2}-r-6=0" .

"\begin{array}{r}

r^{2}-r-6=0 \\

r^{2}-3 r+2 r-6=0 \\

r(r-3)+2(r-3)=0 \\

(r-3)(r+2)=0 \\

r=3,-2

\end{array}"

The roots of the characteristic equation are r_{1}=3, r_{2}=-2.

For n=0 the solution of the recurrence relation is,

"\begin{array}{ll}

a_{0} & =\alpha_{1} r_{1}^{0}+\alpha_{2} r_{2}^{0} \\

3 & =\alpha_{1}(3)^{0}+\alpha_{2}(-2)^{0} \\

3 & =\alpha_{1}+\alpha_{2} \ ...(1)

\end{array} \quad \text { As } a_{0}=3"

For n=1 the solution of the recurrence relation is,

"\begin{array}{ll}

a_{1} & =\alpha_{1} \cdot\left(r_{1}\right)^{1}+\alpha_{2}\left(r_{2}\right)^{1} \\

6 & =\alpha_{1} \cdot(3)^{1}+\alpha_{2}(-2)^{1} \\

6 & =3 \alpha_{1}-2 \alpha_{2}\ ...(2)

\end{array} \quad \text { As } a_{1}=6"

Solve equation (1) and (2) to find the values of "\alpha_{1}, \alpha_{2}" .

Multiply equation (1) by 2 and add equation (1) and (2).

"\begin{aligned}

6 &=2 \alpha_{1}+2 \alpha_{2} \\

6 &=3 \alpha_{1}-2 \alpha_{2} \\

12 &=5 \alpha_{1} \\

\alpha_{1} &=\frac{12}{5}

\end{aligned}"

Thus, the value of "\alpha_{2}" is,

"\begin{array}{r}

\frac{12}{5}+\alpha_{2}=3 \\

\alpha_{2}=\frac{3}{5}

\end{array}"

Therefore, the solution of the recurrence relation is,

"\begin{aligned}

a_{n} &=\alpha_{1} r_{1}^{n}+\alpha_{2} r_{2}^{n} \\

&=\left(\frac{12}{5}\right)(3)^{n}+\left(\frac{3}{5}\right)(-2)^{n}

\end{aligned}"